40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9248     Entry time: Wed Oct 16 19:19:14 2013
Author: Masayuki 
Type: Update 
Category: Green Locking 
Subject: End PDH 60 Hz comb noise in YARM 

[Manasa, Masayuki]

- Motivation

For PRMI + 2arm, we tried to make the ALS control noise better. As this entry we had huge 60 Hz comb noise in PDH loop of YARM.

So we tried to figure out the problem and fix it.

- What we did

We checked which power supply the staff in Y-end are connected to, and change some of them to connect to 1Y4 AC power supply from wall AC. What we changed was
1.Main end laser
2.He-Ne laser
3.Green REFL PD

We checked error signal of PDH control and compared before and after. The 60 Hz peak get better from -80 dBVpk to -90 dBVpk. Also I attached the plot of XARM, privious YARM (the data of Yesterday night), and current YARM ALS in-loop noises. The RMS of ALS in-loop noise of Y-arm get better by factor of 2. However, even the 60 Hz comb noise get better than before, RMS get worse by comb noise. 

We would like to make these noise better at least until these noises don't affect to RMS, so we should continue to check.

Attachment 1: comb_noise.pdf  54 kB  | Hide | Hide all
comb_noise.pdf
ELOG V3.1.3-